A tire 0.450 m in radius rotates at a constant rate of 160 rev/min. Find the speed and acceleration of a small stone lodged in the tread of the tire (on its outer edge).
a)___ m/s

b) What is the acceleration of the stone?___m/s^2

c) tangent to the path or toward the center of the path or away from the center of the path ?
Answer #1

angular speed = 2pi*N / 60= 2pi*160 /60

= 16.75 rad/s

a) v = r*w

= 0.45*16.75 = 7.54 m/s

b) acceleration = r*w^2 = 0.45*16.75^2 = 126.33 m/sw

c) acceleration will be away from the center of the path